OpenStudy (anonymous):

Where does the graph of y = 2x2 − 5x + 3 cross the x-axis?

OpenStudy (anonymous):

it means that find the answer/s of y=0...

OpenStudy (anonymous):

@fatemeh8989 so how do you plug it in and solve

OpenStudy (anonymous):

is that 2x^2?

OpenStudy (anonymous):

@fatemeh8989 yes

OpenStudy (anonymous):

ok... i solve it by delta method: delta = (-5)^2 - 4(2)(3) = 1 x1 = (5+1)/4 = 3/2 x2= (5-1)/4 = 1 need more explanation?